OpenStudy (anonymous):

If s = u + 0.5at2, which equation is solved for t?

OpenStudy (cwrw238):

subtract u from both sides 0.5at^2 = s - u divide both sides by 0.5a: t^2 = 2(s - u) ------ a t = sqrt [2(s - u)/ a]

OpenStudy (anonymous):

thx !!

OpenStudy (anonymous):

thats not a choice :/

OpenStudy (cwrw238):

what are the choices

OpenStudy (anonymous):

t = square root of s minus u over 0.5 times a t = square root of s plus u over 0.5 times a t = square root of s minus u minus 0.5 over a t = square root of s plus u minus 0.5 over a

OpenStudy (cwrw238):

the first one is the same as my answer - dividing by 0.5 is the same as multiplying by 2

OpenStudy (anonymous):

whoops! thanks a lot

OpenStudy (cwrw238):

yw